# Study on physical properties and characteristics of an anisotropic compact star model using Karmarkar Condition in $\mathcal{F}(\mathcal{Q})$ gravity

### v1

@@ IV Conclusion
This study examines a spherically symmetric compact star model with non-perfect fluid. We examined the generic function $\mathcal{F}(\mathcal{Q})$ defined as $\mathcal{F}(\mathcal{Q})=\delta_{1}\mathcal{Q}+\delta_{2}$, where $\delta_{1}$ and $\delta_{2}$ are integration constant. The Karmarkar criteria have been used to develop the differential equation connecting the metric elements of spacetime. By solving this differential equation, we have derived an expression for the metric element $e^{\Upsilon}$ in relation to $e^{\sigma}$. The juxtaposition of the internal spherical metric with the exterior Schwarzschild metric imposed constraints on the arbitrary constants throughout this structure. The results acquired were illustrated visually in Fig. 1– 12. The main results may be described as follows:
@@
  • The compact star’s compactness and mass are shown in Fig. 12. The impact of parameter $\delta_{1}$ on the compactness and mass of the compact star is visible. When the value of $\delta_{1}$ increases, the value of mass and compactness also increases. The values of mass and compactness are zero at the centre of the compact star, and they monotonically increase towards the surface. It is significant to note that there is a mass-radius ratio bound for any stellar structure, which is provided by Buchdahl and is $\frac{2M}{R}\leqslant\frac{8}{9}$ Buchdahl:1959zz . This bound has been computed for the compact star in our model and is displayed in Table 1. This bound is less than $\frac{4}{9}$, demonstrating that our model complies with Buchdahl’s requirements.
  • We have examined the behaviour of various quantities, including metric elements, energy density, radial and tangential pressures, energy and pressure gradients concerning the radial coordinate, anisotropy parameter, $\omega_{r}$ and $\omega_{t}$, energy conditions, adiabatic index, TOV equation, causality conditions, Harrison-Zeldovich-Novikov condition, redshift, mass and compactness for the astronomical object Cen X-3. The results are summarized in Table 2, which confirms the physical plausibility and mathematical coherence of our spherically symmetric model.

### v2

@@ IV Conclusion
This study examines a SS compact star model with non-perfect fluid. We examined the generic function $\mathcal{F}(\mathcal{Q})$ defined as $\mathcal{F}(\mathcal{Q})=\delta_{1}\mathcal{Q}+\delta_{2}$, where $\delta_{1}$ and $\delta_{2}$ are integration constant. The Karmarkar criteria have been used to develop the differential equation connecting the metric elements of spacetime. By solving this differential equation, we have derived an expression for the metric element $e^{\Upsilon}$ in relation to $e^{\sigma}$. The juxtaposition of the internal spherical metric with the exterior Schwarzschild metric imposed constraints on the arbitrary constants throughout this structure. The results acquired were illustrated visually in Fig. 1– 12. The main results may be described as follows:
@@
  • However, Fig. 12 displays the mass and compactness of the compact star. The influence of factor $\delta_{1}$ on the compactness and mass of the star is evident. As the magnitude of $\delta_{1}$ grows, both mass and compactness correspondingly rise. The mass and compactness values are 0 in the center of star and rise gradually approaching its outermost layer. It is important to acknowledge that a mass-radius ratio constraint exists for every star configuration, as established by Buchdahl, which is expressed as $\frac{2M}{R}\leqslant\frac{8}{9}$ Buchdahl:1959zz . The bound has been calculated for our star models and is shown in Table 1. This constraint is inferior than $\frac{4}{9}$, indicating that our model adheres to Buchdahl’s criteria.
  • We have examined the behaviour of various quantities, including metric elements, energy density, radial and tangential pressures, energy and pressure gradients concerning the radial coordinate, anisotropy parameter, $\omega_{r}$ and $\omega_{t}$, energy conditions, adiabatic index, TOV equation, causality conditions, Harrison-Zeldovich-Novikov condition, redshift, mass and compactness for the astronomical object Cen X-3. The results are summarized in Table 2, which confirms the physical plausibility and mathematical coherence of our SS model.